Calculate the change in internal energy (ΔrU) for the following reaction at 1.00 bar and 298 K. Express your answer in kJ/mol. ?8?18(?)+17?2(?) → 8??2(?)+9?2?(?) ∆??=−5430??/??l
Answer #1

Given rH = - 5430 kJ/mol

According to 1st law of thermodynamics:

rH = rU + P*V

Also P*V = - n(g)*RT

For the given reaction,

n(g) = (moles of gaseous products) - (moles of gaseous reactants)

=> n(g) = 8 - 17 = - 9 mol

Now P*V = - (- 9 mol * 8.314 J/mol.K * 298K) = 22298 J/mol * (1kJ / 1000 J) = 22.298 kJ/mol

=> rH = rU + P*V

=> rU = rH - P*V

=> rU = −5430 kJ/mol - 22.298 kJ/mol

=> rU = - 5452.3 kJ/mol (Answer)